# Helix: Distributed Serving of Large Language Models via Max-Flow on Heterogeneous GPUs

### v1

@@
Abstract
This paper introduces Helix, a distributed system for high-throughput, low-latency large language model (LLM) serving on heterogeneous GPU clusters. A key idea behind Helix is to formulate inference computation of LLMs over heterogeneous GPUs and network connections as a max-flow problem for a directed, weighted graph, whose nodes represent GPU instances and edges capture both GPU and network heterogeneity through their capacities. Helix then uses a mixed integer linear programming (MILP) algorithm to discover highly optimized strategies to serve LLMs. This approach allows Helix to jointly optimize model placement and request scheduling, two highly entangled tasks in heterogeneous LLM serving. Our evaluation on several heterogeneous cluster settings ranging from 24 to 42 GPU nodes show that Helix improves serving throughput by up to 2.7$\times$ and reduces prompting and decoding latency by up to 2.8$\times$ and 1.3$\times$, respectively, compared to best existing approaches.
footnotetext: Correspond to: Yixuan Mei <yixuanm@andrew.cmu.edu>
1 Introduction
Generative large language models (LLMs) such as GPT-4 [1] and LLaMA-3 [20] have demonstrated exceptional capabilities of creating natural language texts across a spectrum of application domains, including chatbot [25], coding assistant [19, 30], and task automation [11]. However, the increasingly large model sizes and high computational requirements of modern LLMs make it challenging to serve them cheaply and efficiently on modern cloud platforms. In particular, most of today’s LLM serving systems (e.g., Orca [46] and vLLM [16]) target homogeneous GPU clusters [21], where all GPUs are of the same type and have identical memory capacity and compute resources. Due to the increasing model sizes, serving LLMs using homogeneous GPUs requires an increasing number of GPUs, as shown in Table 1. In addition, serving state-of-the-art LLMs used in industry requires even more resources. We observed it’s nearly infeasible to allocate GPUs of this magnitude within a single cloud region, which is also acknowledged by other recent works [36, 45].
Due to advances in GPU architectural designs and the incremental deployment of them over time, modern cloud platforms increasingly consist of a mix of GPU types. These heterogeneous GPU instances are spread to datacenters around the world and collectively offer significantly larger memory capacity and more compute resources than individual GPU types, enabling a more accessible and scalable approach to LLM serving. Similarly, there is also a trend of using volunteer consumer GPUs to address the GPU scarcity problem [32, 7, 47]. However, in contrast to homogeneous GPU instances, deploying LLMs on geo-distributed heterogeneous instances necessitates accommodating various GPU devices and network conditions.
Table 1: Minimum numbers of GPUs required to serve LLMs in existing homogeneous serving systems. We use half of GPU memory to store model parameters and the other half for key-value cache.
@@ 1 Introduction
LLaMA-3 [41] 70 billion 12 7 4
GPT-3 [1] 175 billion 30 18 9
Grok-1 [44] 314 billion 53 32 16
Prior work has introduced several systems for running machine learning (ML) computation over heterogeneous devices [26, 49] or geo-distributed environments [12, 31]. But most of them are designed for long-running training workloads and cannot adapt to LLM serving scenarios with real-time inference requests. For example, the most closely relevant work to ours is Petals [4], which uses pipeline model parallelism to partition an LLM into multiple stages and employs a greedy algorithm to assign heterogeneous GPUs to these stages in a round-robin fashion. Such throughput-optimized design is effective for long-duration tasks but lacks the responsiveness required for interactive LLM applications like chatbot that demand rapid responses.
To leverage heterogeneous GPU resources for LLM serving, we propose Helix, a distributed system that enables high-throughput, low-latency LLM serving on heterogeneous GPUs. The key idea behind Helix is to formulate the execution of LLM serving over heterogeneous GPUs and networks as the data flow problem under the constraints of diverse GPU computing capabilities, memory capacities, as well as complex inter-GPU connections. To the best of our knowledge, Helix is the first serving system designed for geo-distributed heterogeneous GPUs. Next, we introduce the key challenges Helix must address and Helix’s solutions to them.
First, due to the increasing size of LLMs, serving them on modern GPUs requires employing tensor [35] and pipeline [26, 13] model parallelism to partition an LLM into multiple stages and place these stages on different GPUs, a task we term model placement. Homogeneous serving systems (e.g., Orca [46]) partition an LLM into equal-sized stages and assign them to GPUs. This approach results in suboptimal utilization of high-performance GPUs as it accommodates the memory and computational limitations of less powerful GPUs. Existing heterogeneity-aware serving systems (e.g., Petals [4]) rely on different heuristics to partition a model into stages and assign them to GPUs. Existing heuristics do not simultaneously consider both GPU and network heterogeneity.
Helix formulates model placement as a max-flow problem of a directed, weighted graph, whose nodes represent GPU instances and edges capture both GPU and network heterogeneity through their capacities in the max-flow problem. Helix then uses a mixed integer linear programming (MILP) algorithm to discover highly optimized model placement strategies, which largely outperform the heuristic methods used in prior work [4].
A second challenge Helix must address is request scheduling. To serve an LLM request, Helix needs to select a pipeline of GPU instances to compute all layers of the LLM. Existing systems generally employ a group of fixed pipelines and assign requests to these pipelines in a round-robin fashion. Using fixed pipelines is not flexible enough to accommodate the heterogeneous compute and network conditions and often causes under-utilization. Instead, Helix introduces per-request pipelines, where each request is assigned its own pipeline. As a result, the total number of potential pipelines is equal to the number of paths from source to sink in the graph representation of the cluster, which offers sufficient flexibility for Helix to maximally utilize the full capacity of GPU instances and network connections between them.
@@
We have implemented Helix on top of vLLM [16] and evaluated it on three heterogeneous clusters ranging from 24 to 42 nodes, with up to 7 different node types. The models we evaluated include LLaMA 30B and 70B. Compared to heterogeneity-aware baselines, Helix improves serving throughput by up to 2.7$\times$ while reducing average prompting and decoding latency by up to 2.8$\times$ and 1.3$\times$, respectively.
In summary, this paper makes the following contributions: (1) the first system for high-throughput, low-latency LLM serving on heterogeneous GPUs, (2) a max-flow formulation for LLM serving and an MILP algorithm to optimize model placement; (3) per-request pipelines to maximize GPU utilization; and (4) an implementation of our techniques and an evaluation on various LLM benchmarks. We will release Helix after paper review to facilitate future research.
@@
Figure 1: Examples of sub-optimal model placement and request schedule.  1(a)) all GPUs and network condition in this example. The order of compute capacity is: A100 > L4 > T4;  1(c)) Model placement by uniformly partition the model, then allocate devices by a balanced compute capacity;  1(b)) Co-optimizing model partition and device placement to make the compute capacity more balanced;  1(d)) Co-optimizing model partition, device placement, and request scheduling.
2 Background
2.1 LLM Architecture and Serving
Most of today’s LLMs adopt a decoder-only Transformer architecture [29, 5], which begins by converting a natural language query into a sequence of tokens. The model then converts each token into a hidden state vector, whose size is referred to as the model’s hidden size. A Transformer model comprises of input and output embeddings and a series of identical Transformer layers, each consisting of a self-attention and a feed-forward block. A self-attention block calculates the ‘affinity’ between every pair of tokens and updates each token’s hidden states based on this contextual relevance score. Feed-forward blocks independently modify each token’s hidden state through a non-linear function.
@@ 2.1 LLM Architecture and Serving
To address these challenges, Orca [46] presented iteration-level scheduling, which updates a batch at every iteration to avoid resource retention when a request is completed but others in the same batch need more iterations; vLLM [16] introduces PagedAttention, managing memory for KV-cache with identical pagesand allocating a new page only when a request has used up all its pages; multi-query [33] and group-query attention [3] modifies the self-attention mechanism to reduce the size of KV-cache stored for each token.
2.2 Distributed Model Serving
Open source LLMs now feature up to hundreds of billions of parameters, far exceeding the memory capacity of a single GPU. Consequently, serving an LLM requires multiple GPUs operating in parallel. Tensor Parallelism (TP) [35] partitions the weight of each operator among GPUs, gathering the partial results on each device via an AllReduce/AllGather operation. However, TP is highly sensitive to network conditions. For every Transformer layer, it needs two communications. As a result, TP has a significant overhead in high-latency networks, and is only used among GPUs within a node.
Conversely, Pipeline Parallelism (PP) [13] assigns different operators (typically multiple layers) across GPUs to create multiple pipeline stages. It then splits inputs into microbatches, running them through the pipeline. PP only transmits the activation tensor at the boundary of pipeline stages. Hence, PP is much less network-sensitive. However, it is challenging to perfectly partition both the model and input batch, which results in pipeline bubbles. As a result, PP suffers from the device idle at pipeline bubbles, necessitating careful schedule to be performant [2].
Traditional data center setups typically assume homogeneous clusters: uniform nodes with a uniform bandwidth. As a result, models are evenly partitioned into pipeline stages and assigned to each device. However, with the growing size of LLM to be served and the shortage of the latest generation GPUs, serving LLM with heterogeneous devices emerges as a critical demand, which is not well studied in previous efforts.
Serving LLMs on heterogeneous GPUs needs to consider heterogeneity for both hardware and network. For example, evenly partition of model layers may under-utilizates more capable devices. Additionally, nodes may be located in multiple zones. The bandwidth discrepancies between different zones must be considered.
To tackle these challenges, Petals [4] introduced a greedy device allocation and request routing. It focused on a decentralized setup, which allows hot plug-in and -out of devices. For device placement, it allocates new devices to the least efficient pipeline stages. For network topology, it routes each request by prioritizing the low-latency network. However, such a greedy online approach is sub-optimal in the heterogeneous cluster scenario, where the information of all nodes is a known priori. SWARM [31] focused on model training in a heterogeneous cluster. It evenly partitions the model into pipeline stages. When routing a request to the next pipeline stage, it selects the replica based on the real-time throughput of each candidate.
3 Optimization Formulation in Helix

### v2

# Helix: Serving Large Language Models over Heterogeneous GPUs and Network via Max-Flow

@@
Abstract.
This paper introduces Helix, a distributed system for high-throughput, low-latency large language model (LLM) serving in heterogeneous GPU clusters. The key idea behind Helix is to formulate inference computation of LLMs over heterogeneous GPUs and network connections as a max-flow problem on directed, weighted graphs, whose nodes represent GPU instances and edges capture both GPU and network heterogeneity through their capacities. Helix then uses a mixed integer linear programming (MILP) algorithm to discover highly optimized strategies to serve LLMs on heterogeneous GPUs. This approach allows Helix to jointly optimize model placement and request scheduling, two highly entangled tasks in heterogeneous LLM serving. Our evaluation on several heterogeneous clusters ranging from 24 to 42 GPU nodes shows that Helix improves serving throughput by up to 3.3$\times$ and reduces prompting and decoding latency by up to 66% and 24%, respectively, compared to existing approaches. Helix is available at https://github.com/Thesys-lab/Helix-ASPLOS25.
large language model serving, system for ML, distributed systems, cloud computing
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30–April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30–April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3669940.3707215isbn: 979-8-4007-0698-1/25/03ccs: Computer systems organization Cloud computingccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Parallel computing methodologies
1. Introduction
Table 1. Minimum numbers of GPUs required to serve LLMs in existing homogeneous serving systems. We use half of GPU memory to store model parameters and the other half for key-value cache.
@@ 1. Introduction
LLaMA-2 (touvron2023llama2, ) 70 billion 12 7 4
GPT-3 (achiam2023gpt, ) 175 billion 30 18 9
Grok-1 (xaigrok, ) 314 billion 53 32 16
LLaMA-3 (dubey2024llama, ) 405 billion 68 41 21
Table 2. Availability of different GPU instances in 6 regions on Google Compute Engine (googlecloud2024compute, ).
GPU Type
Region H100 A100 80GB A100 40GB L4 T4 V100
us-central-1 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
us-east-4 $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
us-east-1 $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
eu-west-3 $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
asia-ne-1 $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
asia-ne-3 $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Generative large language models (LLMs) such as GPT-4 (achiam2023gpt, ) and LLaMA-3 (metaIntroducingMeta, ) have demonstrated exceptional capabilities of creating natural language texts across a spectrum of application domains, including chatbot (chatgpt, ), coding assistant (luo2023wizardcoder, ; roziere2023code, ), and task automation (hong2023metagpt, ). However, the increasingly large model sizes and high computational requirements of modern LLMs make it challenging to serve them cheaply and efficiently on modern cloud platforms. In particular, most of today’s LLM serving systems (e.g., Orca (yu2022orca, ) and vLLM (kwon2023vllm, )) target homogeneous GPU clusters (miao2023towards, ), where all GPUs are of the same type and have identical memory capacity and compute resources. Due to increasing model sizes, serving LLMs using homogeneous GPUs requires an increasing number of GPUs, as shown in Table 1. In addition, serving state-of-the-art LLMs used in industry requires even more resources. Recent works have identified that it is increasingly difficult to allocate GPUs of this magnitude within a single cloud region (strati2024ml, ; yang2023skypilot, ).
Due to advances in GPU architectural designs and the incremental deployment of them over time, modern cloud platforms increasingly consist of a mix of GPU types. Table 2 illustrates the heterogeneous GPU deployment in Google Compute Engine  (googlecloud2024compute, ), where datacenters are equipped with various NVIDIA GPUs including H100, A100, V100, L4, and T4. These heterogeneous GPU instances are spread across datacenters around the world and collectively offer significantly larger memory capacity and more compute resources than individual GPU types, enabling a more accessible and scalable approach to LLM serving. As Table 3 shows, eight NVIDIA L4 GPUs can offer comparable FP16 compute performance to a single NVIDIA H100 GPU, while providing greater memory capacity, lower power consumption, and a more cost-effective price point. Moreover, the availability of these GPUs vary significantly across regions. We empirically find that obtaining higher quotas and securing GPU instances is much easier in some regions than others on Google Cloud Engine, and the availability of different GPUs vary a lot (see Table 4).
Geo-distributed LLM serving with heterogeneous GPUs enables the aggregation of available GPUs from multiple regions. This approach not only enhances resource utilization but also minimizes LLM serving costs by strategically leveraging the most cost-effective GPU instances across various geographical locations. Similarly, there is also a trend of using volunteer consumer GPUs to address the GPU scarcity problem (ryabinin2020towards, ; diskin2021distributed, ; yuan2022decentralized, ). However, in contrast to homogeneous GPU instances, deploying LLMs on geo-distributed heterogeneous instances necessitates accommodating various GPU devices and network conditions.
Table 3. Properties of GPUs deployed in today’s data centers. We report SXM version for H100 and A100. Data collected from NVIDIA data sheet (nvidia_h100, ; nvidia_a100, ; nvidia_l4, ; nvidia_t4, ).
FP16 Memory Bandwidth Power Price
GPU (TFLOPs) (GB) (GB/s) (W) (USD)
H100 (nvidia_h100, ) 1979 80 3350 700 25k $\sim$40k
A100 (nvidia_a100, ) 312 40 1555 400 10k $\sim$15k
L4 (nvidia_l4, ) 242 24 300 72 $\sim$3k
T4 (nvidia_t4, ) 65 16 300 70 $\sim$1k
Table 4. GPU quotas and deployment limits in us-east and asia-southeast during our evaluation of Helix. We measure deployment limits at two distinct time periods.
Quota Max Deployed
Region A100 40GB L4 T4 A100 40GB L4 T4
us-east 8 8 16 0 0 16
asia-southeast 8 24 32 4 12 >20
Prior work has introduced several systems for running machine learning computation over heterogeneous devices or geo-distributed environments. However, prior attempts either focus on long-running training workloads (park2020hetpipe, ; zhang2024hap, ; hsieh2017gaia, ; ryabinin2023swarm, ; mei2024realhf, ), which cannot adapt to LLM serving scenarios with real-time inference requests, or focus on decentralized serving with volunteer computing (borzunov2022petals, ; borzunov2024distributed, ), which lack the global coordination necessary to efficiently use GPU and network resources in clusters.
To efficiently serve LLMs over heterogeneous GPUs and network, we propose Helix, a distributed system for high-throughput, low-latency LLM serving. Helix’s key idea is to formulate the execution of LLM serving over heterogeneous GPUs and network as a data flow problem under the constraints of diverse GPU computing capabilities, memory capacities, as well as complex inter-GPU connections. Helix leverages mixed integer linear programming to determine optimal model placement under these constraints. To accommodate heterogeneity, Helix introduces per-request pipelines, where each request has its own independent pipeline for scheduling. This combination of flow-based formulation and per-request pipelines enables Helix to achieve high GPU utilization in heterogeneous and geo-distributed GPU clusters. We will discuss the challenges and Helix’s solutions in Sec. 3.
@@
We have implemented Helix on top of vLLM (kwon2023vllm, ) and evaluated it on three heterogeneous clusters ranging from 24 to 42 nodes, with up to 7 different node types. The models we evaluated include LLaMA-1 30B and LLaMA-2 70B. Compared to heterogeneity-aware baselines, Helix improves serving throughput by up to 3.3$\times$ while reducing average prompting and decoding latency by up to 66% and 24%.
In summary, our contributions are:
  • A system for LLM serving in heterogeneous and geo-distributed GPU clusters.
  • A max-flow formulation for LLM serving and an MILP-based algorithm to optimize model placement.
  • Flexible flow-based per-request pipelines to maximize GPU utilization.
  • An implementation of our techniques and an evaluation on various LLM benchmarks.
2. Background
2.1. LLM Architecture and Serving
Most of today’s LLMs adopt a decoder-only Transformer architecture (radford2019language, ; brown2020language, ), which begins by converting a natural language query into a sequence of tokens. The model then converts each token into a hidden state vector, whose size is referred to as the model’s hidden size. A Transformer model comprises of input and output embeddings and a series of identical Transformer layers, each consisting of a self-attention and a feed-forward block. A self-attention block calculates the ‘affinity’ between every pair of tokens and updates each token’s hidden states based on this contextual relevance score. Feed-forward blocks independently modify each token’s hidden state through a non-linear function.
[…]
To address these challenges, Orca (yu2022orca, ) presented iteration-level scheduling, which updates a batch at every iteration to avoid resource retention when a request is completed but others in the same batch need more iterations; vLLM (kwon2023vllm, ) introduces PagedAttention, managing memory for KV-cache with identical pages and allocating a new page only when a request has used up all its pages; multi-query (shazeer2019fast, ) and group-query attention (ainslie2023gqa, ) modifies the self-attention mechanism to reduce the size of KV-cache stored for each token.
2.2. Distributed Model Serving
Open source LLMs now feature up to hundreds of billions of parameters, far exceeding the memory capacity of a single GPU. Consequently, serving an LLM requires multiple GPUs operating in parallel. Tensor Parallelism (TP) (shoeybi2019megatron, ) partitions the weight of each operator among GPUs, gathering the partial results on each device via an AllReduce/AllGather operation. However, TP is highly sensitive to network conditions. For every Transformer layer, it needs two communications. As a result, TP has a significant overhead in high-latency networks, and is only used among GPUs within a node.
Conversely, Pipeline Parallelism (PP) (huang2019gpipe, ) assigns different operators (typically multiple layers) across GPUs to create multiple pipeline stages. It then splits inputs into microbatches, running them through the pipeline. PP only transmits the activation tensor at the boundary of pipeline stages. Hence, PP is much less network-sensitive. However, it is challenging to perfectly partition both the model and input batch, which results in pipeline bubbles. As a result, PP suffers from the device idle at pipeline bubbles, and necessitating careful schedule to be performant (agrawal2023sarathi, ).
Traditional data center setups typically assume homogeneous clusters: uniform nodes with a uniform bandwidth. As a result, models are evenly partitioned into pipeline stages and assigned to each devices. As LLMs grow in size and the latest generation GPUs remain scarce, deploying these models across heterogeneous computing devices has become a critical necessity, a challenge that previous research has not adequately addressed.
Deploying LLMs across geo-distributed, heterogeneous GPU clusters requires careful consideration of both hardware capabilities and network characteristics. Simple equal distribution of model layers across devices fails to maximize the potential of more powerful hardware. Moreover, the significant bandwidth differences between intra- and inter-regional network connections must inform both model placement and request scheduling decisions, particularly when infrastructure spans multiple geographical regions.
No prior work has focused on LLM serving on heterogeneous GPU clusters. The most related work is Petals (borzunov2022petals, ), which performs decentralized LLM serving with volunteer computing. Users contribute GPUs to form a decentralized swarm and newly joined machines greedily serve the pipeline stages with least compute capacity. For request scheduling, users greedily choose the server with lowest latency to them. Petals is effective for volunteer computing, but the lack of global coordination makes it unable to fully utilize the GPUs and network when the cluster is a known priori. Another relevant work is SWARM (ryabinin2023swarm, ), which performs DNN training in heterogeneous clusters. It evenly partitions the model into pipeline stages. When routing a request to the next pipeline stage, it selects the replica based on real-time throughput of candidates. Our evaluation shows that such simple heuristics are not enough to achieve good performance in geo-distributed clusters with heterogeneous GPUs and network.
@@ 2.2. Distributed Model Serving
Figure 1. Examples of sub-optimal model placement and request schedule.  1(a)) all GPUs and network condition in this example. The order of compute capacity is: A100 ¿ L4 ¿ T4;  1(b)) Model placement by uniformly partition the model, then allocate devices by a balanced compute capacity;  1(c)) Co-optimizing model partition and device placement to make the compute capacity more balanced;  1(d)) Co-optimizing model partition, device placement, and request scheduling in a network-aware way.
@@
Figure for model placement.
3. Opportunities and Challenges
Using heterogeneous and geo-distributed GPUs presents new opportunities for LLM serving. As shown in Table 3, multiple commodity GPUs (L4, T4) can match the compute capacity of high-end GPUs (H100, A100) while offering advantages in memory capacity, energy efficiency, and cost. Furthermore, Tables 4 and 7 demonstrate that GPU availability varies significantly across regions, yet inter-region network conditions remain suitable for LLM serving, motivating a geo-distributed approach. However, leveraging heterogeneous and geo-distributed GPUs poses several key challenges.
3.1. Challenge 1: Model Placement
Due to the increasing size of LLMs, serving them on modern GPUs requires employing tensor (shoeybi2019megatron, ) and pipeline (park2020hetpipe, ; huang2019gpipe, ) model parallelism to partition an LLM into stages and place those stages on different GPUs, a task we term model placement. Homogeneous serving systems (e.g., Orca (yu2022orca, )) partition an LLM into equal-sized stages and assign them to GPUs. This approach results in sub-optimal utilization of high-performance GPUs as it accommodates the memory and computational limitations of less powerful GPUs. Existing heterogeneity-aware serving systems (e.g., Petals (borzunov2022petals, )) rely on different heuristics to partition a model into stages and assign them to GPUs. Existing heuristics do not simultaneously consider both GPU and network heterogeneity.
Helix’s solution:
Helix exploits the flexibility of token-level scheduling in LLM serving and formulates model placement as a max-flow problem of a directed, weighted graph, whose nodes represent GPU instances and edges capture both GPU and network heterogeneity through their capacities in the max-flow problem. Helix then uses a mixed integer linear programming (MILP) algorithm to discover highly optimized model placement strategies, which largely outperform the heuristic methods used in prior work (borzunov2022petals, ; ryabinin2023swarm, ). Leveraging the data dependencies and homogeneity of LLM layers, Helix expresses the MILP problem with linear number of variables and constraints relative to the number of compute nodes and network connections, resulting in a tractable problem size.
3.2. Challenge 2: Request Scheduling
A second challenge Helix must address is request scheduling. To serve an LLM request, Helix needs to select a pipeline of GPU instances to compute all layers of the LLM. Existing systems generally employ a group of fixed pipelines and assign requests to these pipelines in a round-robin fashion. Using fixed pipelines is not flexible enough to accommodate the heterogeneous compute and network conditions and often causes under-utilization.
Helix’s solution:
Helix introduces per-request pipelines, where each request is assigned its own pipeline. As a result, the total number of potential pipelines is equal to the number of paths from source to sink in the graph representation of the cluster, which offers sufficient flexibility for Helix to maximally utilize the full capacity of GPU instances and network connections between them.
(a) A 3-node cluster with model placement. Network connections between the coordinator and compute nodes transmit tokens (4 Byte) while others transmit intermediate activations (16 KB)
(b) Graph abstraction of the cluster.
Figure 2. Graph abstraction of a 3-node cluster with given model placement. Numbers on the edges in Fig. 2(b) represent their capacity, which is the number of tokens that can pass through the edges per second. Max flow between source and sink equals the max serving throughput of the cluster.
[…]
Figure for network flow.
4. Optimization Formulation in Helix